OpenStudy (anonymous):

Help Please (medals) The volume of a cube is found by using the formula l^3, where l is the side length. If the side length is 4x^3, what is the volume of the cube? A.64x^6 B.12x^9 C.4x^6 D.64x^9

jigglypuff314 (jigglypuff314):

so V = L^3 and if L = (4x^3) then V = (4x^3)^3 = ?

OpenStudy (anonymous):

uhh... hmm...

OpenStudy (anonymous):

give me a sec

jigglypuff314 (jigglypuff314):

hint: the rule that goes along with this would be (x^m)^n = x^(m*n)

OpenStudy (anonymous):

oh dam i see thank you jiggly puff :D

jigglypuff314 (jigglypuff314):

I'm glad I could help :)